OpenStudy (anonymous):

find D.E x^2/a^2+y^2/b^2=1 ans. xyy''+xy'^2=yy'

OpenStudy (anonymous):

diff. twice and eliminate a^2 and b^2 from three equations.

OpenStudy (anonymous):

how @surjithayer

OpenStudy (anonymous):

help please

ganeshie8 (ganeshie8):

Have you used the surjithayer's hint and differentiated twice ?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

show you work please

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

number 14

ganeshie8 (ganeshie8):

whats the derivative of \(\dfrac{x^2}{a^2}\) with respect to \(x\) ?

OpenStudy (anonymous):

but i think , may solution is not right

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

Keep in mind, \(a\) and \(b\) are constants here

ganeshie8 (ganeshie8):

\[\dfrac{x^2}{a^2}+\dfrac{y^2}{b^2}=1\tag{1}\] differentiating implicitly with respect to \(x\) gives \[\dfrac{2x}{a^2} + \dfrac{2yy'}{b^2} = 0\tag{2}\] yes ?

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

but , i did it first , but my answer was not right

ganeshie8 (ganeshie8):

your work in that picture is wrong you're not treating a^2 as constant, instead trying to differentiate it and writing 2a.. which is wrong

OpenStudy (anonymous):

can you help please

ganeshie8 (ganeshie8):

\[\dfrac{x^2}{a^2}+\dfrac{y^2}{b^2}=1\tag{1}\] differentiating implicitly with respect to \(x\) gives \[\dfrac{2x}{a^2} + \dfrac{2yy'}{b^2} = 0 \implies a^2 = -\dfrac{b^2x}{yy'} \] Plug this in first equation and get \[\dfrac{x^2}{-b^2x/(yy')}+\dfrac{y^2}{b^2}=1 \] \[\implies -xyy' +y^2 = b^2 \tag{2}\] differentiate this again and you're done!

OpenStudy (anonymous):

@ganeshie8 gets :)) thank you very much :))

ganeshie8 (ganeshie8):

yw, are you able to differentiate it again and arrive at the answer ?

OpenStudy (anonymous):

@ganeshie8 im very confused

ganeshie8 (ganeshie8):

remember product rule ?

OpenStudy (anonymous):

yeah i remember

OpenStudy (anonymous):

how about the b^2 ?

ganeshie8 (ganeshie8):

b^2 is just a constant, like, 3^2 or 5^2

ganeshie8 (ganeshie8):

it doesn't change, so the derivative is 0

OpenStudy (anonymous):

\[-x y y'' . -1 + x y^2.y'=b^2\]

ganeshie8 (ganeshie8):

\[\dfrac{x^2}{a^2}+\dfrac{y^2}{b^2}=1\tag{1}\] differentiating implicitly with respect to \(x\) gives \[\dfrac{2x}{a^2} + \dfrac{2yy'}{b^2} = 0 \implies a^2 = -\dfrac{b^2x}{yy'} \] Plug this in first equation and get \[\dfrac{x^2}{-b^2x/(yy')}+\dfrac{y^2}{b^2}=1 \] \[\implies -xyy' +y^2 = b^2 \tag{2}\] differentiating w.r.t \(x\) again \[-(yy' +y'xy' + y''xy) + 2yy'= 0\] \[yy' = xyy'' + x{y'}^2\]

ganeshie8 (ganeshie8):

Recall below product rules \[\left(f(x)*g(x)\right)' = f'(x)g(x) + f(x)g'(x)\] \[\left(f(x)*g(x)*h(x)\right)' = f'(x)g(x)h(x) + f(x)g'(x)h(x) + f(x)g(x)h'(x)\]

OpenStudy (anonymous):

@ganeshie8 i got the answer :))))

OpenStudy (anonymous):

how about x^2+axy+bx+c=0

ganeshie8 (ganeshie8):

Give it a try first

OpenStudy (anonymous):

i have a answer

OpenStudy (anonymous):

@ganeshie8 number 11

ganeshie8 (ganeshie8):

Looks you forgot to use product rule here ? 

OpenStudy (usukidoll):

isn't that we have to find the second and third derivatives and then plug them back into the equation to check if the equation we are given (that ans. thing) holds true?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (usukidoll):

x and y's are variables treat a b c as constants which when the derivative is taken it goes to 0

OpenStudy (anonymous):

ai derivative of x is 1 = a (1) ?

ganeshie8 (ganeshie8):

@UsukiDoll this is like constructing the differential equation from the given general solution by eliminating the arbitrary constants

OpenStudy (usukidoll):

oh I see so we're not verifying... we're finding a DE

ganeshie8 (ganeshie8):

yeah we can think of the equations of given ellipses/parabolas as general solutions

OpenStudy (anonymous):

yeah , were finding D.E

ganeshie8 (ganeshie8):

@jacalneaila did you find ur mistake in #11 ?

OpenStudy (anonymous):

yeah, i try to solve ut again , but may answer is different now :(

OpenStudy (anonymous):

*it

ganeshie8 (ganeshie8):

show ur work

OpenStudy (anonymous):

wait , i dont have a cellphone

ganeshie8 (ganeshie8):

#11 should be fairly simple compared to the earlier one

OpenStudy (anonymous):

\[ax+yy'+a(1)? \]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):